# Solving Optimal Experiment Design with Mixed-Integer Convex Methods

### v6

@@ 5 Computational experiments
Note that both the objectives (D-Opt) and (A-Opt) are only well defined if the information matrix $X(\mathbf{x})$ has full rank. This is the case for the Fusion Problem, not necessarily for the Optimal Problem. Both Boscia.jl and Co-BnB require a feasible starting point $\mathbf{z}_{0}$. For its construction, we find a set $S\subset[m]$ of $n$ linearly independent experiments, i.e. $n$ linearly independent rows of $A$. Assign those experiments their upper bound. If the sum $\sum\mathbf{z}_{0}$ exceeds $N$, remove 1 from the experiment with the largest entry. If the sum is less than $N$, pick an experiment in $[m]\backslash S$ at random and assign it as many runs as possible. Repeat until the sum is equal to $N$. Note that due to the monotonic progress of both first-order methods, the current iterate will never become domain infeasible, i.e. singular.
@@ Results.
An overview of the results of the computational experiments is given in Table 2 and extensive results can be found in Section D. The new framework Boscia.jl shows a superior performance compared to the other methods both in terms of time and number of instances solved to optimality. In comparison to the Outer Approximation methods, for example, it solves nearly twice as many instances.
Co-BnB is the only competitive method. It mostly solves fewer instances than Boscia.jl, the exception being the A-Fusion Problem with both data sets. With independent data, Co-BnB is faster solving the same number of instances. In case of the correlated data set, Co-BnB solves roughly twice as many instances as Boscia.jl. In Table 5, we see that Boscia.jl and Co-BnB have similar magnitudes in number of nodes for both A-Fusion Problems whereas for all other problems Boscia.jl has significantly fewer nodes than Co-BnB. As stated previously, node evaluation in Boscia.jl is more computationally expensive. The fact that the magnitudes of nodes is similar points to the optimal solution being strictly in the interior. Note that in general, Co-BnB fares well for the instances where $n=\lfloor m/10\rfloor$ as it was designed with such problems. It struggles for the instances where $n=\lfloor m/4\rfloor$. In Figure 3, we see that the absolute gap of Co-BnB increasing drastically as the fraction between number of parameters and number experiments grow larger. While Boscia.jl also shows an increase, it is much more moderate. Co-BnB implements no advanced Branch-and-Bound tree specifications like a better traverse strategy or branching strategy. A greater value of $N$ naturally increases the size of the tree and the number of nodes to be processed. Boscia.jl has the advantage here since it finds many integer feasible points while solving the relaxations which have the potential to improve the incumbent. A better incumbent, in turn, lets us prune non-improving nodes early on. Overall, fewer nodes have to be investigated by Boscia.jl.
@@
In terms of time, Boscia.jl also shows promising results, especially for instances of larger scale. For small-scale instances, the Outer Approximation approaches are fast, in some cases faster than Boscia.jl. This can be observed in Figures 4 and 5. The plots also highlight that the Outer Approximation Schemes solve fewer instances than the two Branch-and-Bound Approaches. It should be noted that SCIP + OA simply timed out often. On the other hand, the MIP solvers for the conic formulations with Pajarito.jl often struggled to solve the resulting MIP problems. SOCP, for example, only solved about a quarter of the instances of the D-Fusion Problem and D-Optimal Problem. In addition, the many solutions provided by SOCP in case of a timeout were not always domain feasible, and thus, the original objective function could not be evaluated at the provided solutions. The superiority of the Branch-and-Bound approaches stems from the fact that the relaxations of both of them keep the simple feasible region intact. Whereas the Outer Approximation schemes add many additional constraints, i.e. cuts, resulting in larger LPs to be solved, see Table 5 for the average number of cuts. Furthermore, these cuts are dense, i.e. have many nonzero entries, leading to an increased difficulty for the LP solvers.
@@
For the (GTI-Opt) Problem, the results are given in Table 3. Note that we used Boscia.jl on both the (GTI-Opt) and the (logGTI-Opt), the latter titled as "Boscia Log" in the table. Boscia.jl also shows a better performance here than Co-BnB and SCIP + OA. Interestingly, the (logGTI-Opt) objective seems to lead to a better conditioning of the problem. This further motivates an investigation into generalized self-concordance of those objectives.
In Table 4, we see the relative and absolute gap of Boscia.jl and Co-BnB for instances with sizes from 300 and 500 variables. Boscia.jl showcases better gap values especially for the problems under the D-criterion. Note that almost none of the instances were solved to optimality.
We performed a hyperparameter search for Boscia.jl the results of which can be found in Tables 8, 9 and 10.
@@
It can be observed in Table 2 that most solvers solve fewer instances under the A-criterion. The notable exception is SOCP on the A-Fusion Problem where it solves more instances to optimality compared to the D-criterion.
Taking a look at some example contour plots shown in Figure 6, we observe that the contour lines for the A-criterion are steeper than those of the D-criterion for both the Optimal Problem and the Fusion Problem, respectively. This points to the conditioning increasing.
@@
Interestingly, the independent data leads to proof of optimality, i.e. the optimal solution is found early on and the lower bound has to close the gap, regardless of the problem, see Figures 7(a) and 7(c). The difference in problems has, however, an impact on how fast the lower bound can catch up with the incumbent. In the case of the Optimal Problem there is likely a larger region around the optimal solution where the corresponding points/designs $\mathbf{x}$ provide roughly the same information. These other candidates have to be checked to ensure the optimality of the incumbent and thus the solving process slows down. Note that Boscia.jl can utilize strong convexity and sharpness of the objective to improve the lower bound. However, the improvement via sharpness is currently only applicable if the optimum is unique which is not the case.
On the other hand, the correlated data leads to solution processes that are very incumbent-driven, i.e. most improvement on the dual gap stems from the improving incumbent, not from the lower bound, as seen in Figures 7(b) and 7(d). Incumbent-driven solution processes can be identified by the dual gap making sudden jumps and the absence of (a lot of) progress between these jumps. As before, the solution process speed depends on the Problem. In Figure 7(b), the dual gap makes big jumps throughout most of the solving process, in contrast to the dual gap Figure 7(d). This indicates that the optimal solution of the Fusion Problem is strictly in the interior and thus, will not be found early as a vertex of a relaxation. The key ingredient for improvement in this case will be the incorporation of more sophisticated primal heuristics in Boscia.jl.
@@
For more extensive results like number of cuts/nodes, see Table 5 in Section D.
@@
For extensive results, see Tables 6 and 7 in Section D.
@@ A.1 General power of a matrix
Combining Lemma A.17 and Lemma A.18, we get the following corollary.
Corollary A.19.
@@ A.2 Convexity of the objectives
Observe that $I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}$ is a positive definite matrix and its eigenvalues are $1+t\lambda_{i}$ with $\lambda_{i}$ being the $i$-th eigenvalue of $A^{-\frac{1}{2}}BA^{-\frac{1}{2}}$. Note that this also means $01+t\lambda_{i}>$ 0.
@@
Using Lemma A.20 for the derivative yields
@@
By Corollary A.19, $A^{-p}$ is positive definite and thus, has a unique root $A^{-p/2}$. Using the traces cyclic property, we find
@@ A.3 Additional smoothness proofs
Let $A\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Then we can define $h(t)=-\log\det\left(A+tB\right)$. By Lemma A.22, $f(X)$ is convex and the second derivative of $h$ is given by
@@
By using a similar idea to Lemma A.29, we find
@@
Let $A\in D$ and $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Define $h(t)=\operatorname{Tr}((A+tB)^{-p})$. In Lemma A.23 the proof for convexity of $g(X)$ is presented. There one can also find the calculation for $h^{\prime\prime}(t)_{\restriction t=0}$.
@@
Since $A^{-p}$ is positive definite by Corollary A.19 and the trace is cyclic, we have
@@
Note that $BB$ is postive semi-definite. Thus, we can use the bound in Fang et al. (1994).
@@
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$. By Lemma A.24, we have
@@
Note that $BB$ is PSD. The proof of Lemma A.29 can be adapted for one of matrices being symmetric and the other positive definite.
@@
In our case, the minimum eigenvalue is bounded around the optimum. However, it is not bounded in general for our feasible region unless we can introduce domain cuts. Hence, Lemmas A.27, A.25 and A.25 are not straight away useful to us.
@@ A.4 Results on positive semi-definite matrices
Lemma A.28 (Courant-Fischer min-max Theorem).
@@
$\geq\lambda_{1}(B)+\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
@@
$\geq\lambda_{1}(B)\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
@@ A.5 Notes on the Frobenius norm
Note that both $A^{\intercal}A$ and $B^{\intercal}B$ are positve semi-definite. Thus, we can use the result in Fang et al. (1994):
@@ B Custom branch-and-bound with coordinate descent for OEDP
In this section, we showcase the changes we made to the solver set-up in Ahipaşaoğlu (2021). Firstly, we state and solve the problems of interest as minimization problems. Secondly, we forgo the $\mathbf{w}N\in{\mathbb{Z}}\,\forall i\in[m]$ in favor of directly having the variables $x$ to encode if and how often the experiments should be run. The D-Optimal Problem is then stated as
@@
compare Ahipaşaoğlu (2021, Section 4.2). The constant $A$, $B$, $C$ and $D$ are defined as
@@
$\theta$ =$=\begin{cases}\min\left\{-\frac{B+\sqrt{B^{2}-A\Delta}}{\Delta},u_{j}-x_{j},x_{k}-l_{k}\right\}&\text{if }\Delta\neq 0\\ \min\left\{-\frac{A}{2B},u_{j}-x_{j},x_{k}-l_{k}\right\}&\text{if }\Delta=0\text{ and }B\neq 0\\ \min\{u_{j}-x_{j},x_{k}-l_{k}\}&\text{if }\Delta=B=0\text{ and }A>0\end{cases}$
@@ C BPCG convergence under sharpness conditions
In this appendix, we prove linear convergence of BPCG on polytopes when the objective function is sharp or presents a Hölderian error bound. For the definition of sharpness, see Definition 3.11. The following proofs rely on a geometric constant of a polytope $\mathcal{P}$.
Definition C.31 (Pyramidal Witdh Braun et al. (2022)).
@@
Theorem C.32 (Linear convergence of BPCG).
@@
We will denote $h_{t}=f(\mathbf{x}_{t})-f^{*}$. The proof adapts that of Tsuji et al. (2021, Theorem 3.2) to the case of a sharp objective. Tsuji et al. (2021, Lemma 3.5) establishes that at any step $t$, the following holds:
@@
If $t$ follows Case (a) of Lemma C.34, then
@@
[…]
If $t$ follows Case (b) of Lemma C.34, then:
@@
Denoting with $\mathcal{T}_{\mathrm{drop}}\subseteq[T]$ the set of drop steps, we can already observe that the sequence $\{h_{t}\}_{t\in[T]\backslash\mathcal{T}_{\mathrm{drop}}}$ follows the hypotheses of Lemma C.35, with parameters $c_{1},c_{2},\alpha$ defined in the statement of the theorem.
We can then observe that the number of drop steps is bounded by the number of FW steps on the sequence. This implies in the worst case that in $12t,t>$ 1 iterations, we can reach the primal gap of $t$ FW and pairwise steps combined. Let us denote $\tau$ the number of steps that were not drop steps up to iteration $t$, then:
@@
Lemma C.33 (Geometric sharpness Braun et al. (2022), Lemma 3.32).
@@
Lemma C.34 (Minimum decrease Tsuji et al. (2021), Lemma 3.4).
@@
Lemma C.35 (Convergence rates from contraction inequalities, Braun et al. (2022), Lemma 2.21).
@@
Finally, we can exploit the recent results of Zhao (2025) on AFW for log-homogeneous self-concordant barrier functions to show that on D-Opt Problem, BPCG converges at a linear rate in primal and FW gaps.
Proposition C.36 (Linear convergence of BPCG on D-Opt (Zhao, 2025)).
@@
The core observation required to extend Zhao (2025, Theorem 4.1) is that the gap produced by BPCG upper-bounds the one produced by AFW. Let $\mathbf{d}^{\text{BPCG}},\mathbf{d}^{\text{AFW}}$ be the search directions produced by BPCG, AFW respectively. A drop step can occur in both algorithms in at most half of the iterations, since each drop step reduces the size of the active set by one. We can focus on “effective” steps in which no vertex is dropped. BPCG either takes a local pairwise step with direction:
@@
and $\mathcal{S}_{t}$ the local active set. Using the same notation, $\mathbf{d}^{\text{AFW}}=\mathbf{a}-\mathbf{x}$. The quantity determining whether a local pairwise step is selected is $\left\langle\nabla f(\mathbf{x}),\mathbf{d}^{\text{BPCG}}\right\rangle$, which needs to be greater than the FW gap $\left\langle\nabla f(\mathbf{x}),\mathbf{x}-\mathbf{v}\right\rangle$ at the current point. Regardless of the direction taken by BPCG, Tsuji et al. (2021, Lemma 3.5) provides:
@@
which is precisely the only inequality required from effective AFW steps. The rest of the proofs from Zhao (2025, Theorem 4.1, Theorem 4.2) follow without relying on the AFW descent direction specifically. ∎
Proposition C.36 unfortunately relies on the fact that the D-Opt objective is a log-barrier in addition to being self-concordant. Since this is not the case for the trace inverse function, we cannot transpose this result to the A-Opt Problem.

### v7

@@ 5 Computational experiments
Note that both the objectives (D-Opt) and (A-Opt) are only well defined if the information matrix $X(\mathbf{x})$ has full rank. This is the case for the Fusion Problem, not necessarily for the Optimal Problem. Both Boscia.jl and Co-BnB require a feasible starting point $\mathbf{z}_{0}$. For its construction, we find a set $S\subset[m]$ of $n$ linearly independent experiments, i.e. $n$ linearly independent rows of $A$. Assign those experiments their upper bound. If the sum $\sum\mathbf{z}_{0}$ exceeds $N$, remove 1 from the experiment with the largest entry. If the sum is less than $N$, pick an experiment in $\S[m]\backslash$ S at random and assign it as many runs as possible. Repeat until the sum is equal to $N$. Note that due to the monotonic progress of both first-order methods, the current iterate will never become domain infeasible, i.e. singular.
@@ Results.
An overview of the results of the computational experiments is given in Table˜2 and extensive results can be found in Section˜D. The new framework Boscia.jl shows a superior performance compared to the other methods both in terms of time and number of instances solved to optimality. In comparison to the Outer Approximation methods, for example, it solves nearly twice as many instances.
Co-BnB is the only competitive method. It mostly solves fewer instances than Boscia.jl, the exception being the A-Fusion Problem with both data sets. With independent data, Co-BnB is faster solving the same number of instances. In case of the correlated data set, Co-BnB solves roughly twice as many instances as Boscia.jl. In Table˜5, we see that Boscia.jl and Co-BnB have similar magnitudes in number of nodes for both A-Fusion Problems whereas for all other problems Boscia.jl has significantly fewer nodes than Co-BnB. As stated previously, node evaluation in Boscia.jl is more computationally expensive. The fact that the magnitudes of nodes is similar points to the optimal solution being strictly in the interior. Note that in general, Co-BnB fares well for the instances where $n=\lfloor m/10\rfloor$ as it was designed with such problems. It struggles for the instances where $n=\lfloor m/4\rfloor$. In Figure˜3, we see that the absolute gap of Co-BnB increasing drastically as the fraction between number of parameters and number experiments grow larger. While Boscia.jl also shows an increase, it is much more moderate. Co-BnB implements no advanced Branch-and-Bound tree specifications like a better traverse strategy or branching strategy. A greater value of $N$ naturally increases the size of the tree and the number of nodes to be processed. Boscia.jl has the advantage here since it finds many integer feasible points while solving the relaxations which have the potential to improve the incumbent. A better incumbent, in turn, lets us prune non-improving nodes early on. Overall, fewer nodes have to be investigated by Boscia.jl.
@@
In terms of time, Boscia.jl also shows promising results, especially for instances of larger scale. For small-scale instances, the Outer Approximation approaches are fast, in some cases faster than Boscia.jl. This can be observed in Figures˜4 and 5. The plots also highlight that the Outer Approximation Schemes solve fewer instances than the two Branch-and-Bound Approaches. It should be noted that SCIP + OA simply timed out often. On the other hand, the MIP solvers for the conic formulations with Pajarito.jl often struggled to solve the resulting MIP problems. SOCP, for example, only solved about a quarter of the instances of the D-Fusion Problem and D-Optimal Problem. In addition, the many solutions provided by SOCP in case of a timeout were not always domain feasible, and thus, the original objective function could not be evaluated at the provided solutions. The superiority of the Branch-and-Bound approaches stems from the fact that the relaxations of both of them keep the simple feasible region intact. Whereas the Outer Approximation schemes add many additional constraints, i.e. cuts, resulting in larger LPs to be solved, see Table˜5 for the average number of cuts. Furthermore, these cuts are dense, i.e. have many nonzero entries, leading to an increased difficulty for the LP solvers.
@@
For the (GTI-Opt) Problem, the results are given in Table˜3. Note that we used Boscia.jl on both the (GTI-Opt) and the (logGTI-Opt), the latter titled as "Boscia Log" in the table. Boscia.jl also shows a better performance here than Co-BnB and SCIP + OA. Interestingly, the (logGTI-Opt) objective seems to lead to a better conditioning of the problem. This further motivates an investigation into generalized self-concordance of those objectives.
In Table˜4, we see the relative and absolute gap of Boscia.jl and Co-BnB for instances with sizes from 300 and 500 variables. Boscia.jl showcases better gap values especially for the problems under the D-criterion. Note that almost none of the instances were solved to optimality.
We performed a hyperparameter search for Boscia.jl the results of which can be found in Tables˜8, 9 and 10.
@@
It can be observed in Table˜2 that most solvers solve fewer instances under the A-criterion. The notable exception is SOCP on the A-Fusion Problem where it solves more instances to optimality compared to the D-criterion.
Taking a look at some example contour plots shown in Figure˜6, we observe that the contour lines for the A-criterion are steeper than those of the D-criterion for both the Optimal Problem and the Fusion Problem, respectively. This points to the conditioning increasing.
@@
Interestingly, the independent data leads to proof of optimality, i.e. the optimal solution is found early on and the lower bound has to close the gap, regardless of the problem, see Figures˜7(a) and 7(c). The difference in problems has, however, an impact on how fast the lower bound can catch up with the incumbent. In the case of the Optimal Problem there is likely a larger region around the optimal solution where the corresponding points/designs $\mathbf{x}$ provide roughly the same information. These other candidates have to be checked to ensure the optimality of the incumbent and thus the solving process slows down. Note that Boscia.jl can utilize strong convexity and sharpness of the objective to improve the lower bound. However, the improvement via sharpness is currently only applicable if the optimum is unique which is not the case.
On the other hand, the correlated data leads to solution processes that are very incumbent-driven, i.e. most improvement on the dual gap stems from the improving incumbent, not from the lower bound, as seen in Figures˜7(b) and 7(d). Incumbent-driven solution processes can be identified by the dual gap making sudden jumps and the absence of (a lot of) progress between these jumps. As before, the solution process speed depends on the Problem. In Figure˜7(b), the dual gap makes big jumps throughout most of the solving process, in contrast to the dual gap Figure˜7(d). This indicates that the optimal solution of the Fusion Problem is strictly in the interior and thus, will not be found early as a vertex of a relaxation. The key ingredient for improvement in this case will be the incorporation of more sophisticated primal heuristics in Boscia.jl.
@@
For more extensive results like number of cuts/nodes, see Table˜5 in Section˜D.
@@
For extensive results, see Tables˜6 and 7 in Section˜D.
@@ A.1 General power of a matrix
Combining Lemma˜5 and Lemma˜6, we get the following corollary.
Corollary 2.
@@ A.2 Convexity of the objectives
Observe that $I+tA^{-\frac{1}{2}}BA^{-\frac{1}{2}}$ is a positive definite matrix and its eigenvalues are $1+t\lambda_{i}$ with $\lambda_{i}$ being the $i$-th eigenvalue of $A^{-\frac{1}{2}}BA^{-\frac{1}{2}}$. Note that this also means $1+t\lambda_{i}>0$.
@@
Using Lemma˜7 for the derivative yields
@@
By Corollary˜2, $A^{-p}$ is positive definite and thus, has a unique root $A^{-p/2}$. Using the traces cyclic property, we find
@@ A.3 Additional smoothness proofs
Let $A\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Then we can define $h(t)=-\log\det\left(A+tB\right)$. By Lemma˜9, $f(X)$ is convex and the second derivative of $h$ is given by
@@
By using a similar idea to Lemma˜16, we find
@@
Let $A\in D$ and $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Define $h(t)=\operatorname{Tr}((A+tB)^{-p})$. In Lemma˜10 the proof for convexity of $g(X)$ is presented. There one can also find the calculation for $h^{\prime\prime}(t)_{\restriction t=0}$.
@@
Since $A^{-p}$ is positive definite by Corollary˜2 and the trace is cyclic, we have
@@
Note that $BB$ is postive semi-definite. Thus, we can use the bound in fang1994inequalities.
@@
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$. By Lemma˜11, we have
@@
Note that $BB$ is PSD. The proof of Lemma˜16 can be adapted for one of matrices being symmetric and the other positive definite.
@@
In our case, the minimum eigenvalue is bounded around the optimum. However, it is not bounded in general for our feasible region unless we can introduce domain cuts. Hence, Lemmas˜14, 12 and 12 are not straight away useful to us.
@@ A.4 Results on positive semi-definite matrices
Lemma 15 (Courant-Fischer min-max Theorem).
@@
[…]
@@
[…]
@@ A.5 Notes on the Frobenius norm
Note that both $A^{\intercal}A$ and $B^{\intercal}B$ are positve semi-definite. Thus, we can use the result in fang1994inequalities:
@@ B Custom branch-and-bound with coordinate descent for OEDP
In this section, we showcase the changes we made to the solver set-up in ahipacsaouglu2021branch. Firstly, we state and solve the problems of interest as minimization problems. Secondly, we forgo the $\mathbf{w}N\in{\mathbb{Z}}\,\forall i\in[m]$ in favor of directly having the variables $x$ to encode if and how often the experiments should be run. The D-Optimal Problem is then stated as
@@
compare ahipacsaouglu2021branch. The constant $A$, $B$, $C$ and $D$ are defined as
@@
$\theta$ $=\begin{cases}\min\left\{-\frac{B+\sqrt{B^{2}-A\Delta}}{\Delta},u_{j}-x_{j},x_{k}-l_{k}\right\}&\text{if }\Delta\neq 0\\ \min\left\{-\frac{A}{2B},u_{j}-x_{j},x_{k}-l_{k}\right\}&\text{if }\Delta=0\text{ and }B\neq 0\\ \min\{u_{j}-x_{j},x_{k}-l_{k}\}&\text{if }\Delta=B=0\text{ and }A>0\end{cases}$
@@ C BPCG convergence under sharpness conditions
In this appendix, we prove linear convergence of BPCG on polytopes when the objective function is sharp or presents a Hölderian error bound. For the definition of sharpness, see Definition˜5. The following proofs rely on a geometric constant of a polytope $\mathcal{P}$.
Definition 7 (Pyramidal Witdh braun2022conditional).
@@
Theorem 6 (Linear convergence of BPCG).
@@
We will denote $h_{t}=f(\mathbf{x}_{t})-f^{*}$. The proof adapts that of tsuji2021sparser to the case of a sharp objective. tsuji2021sparser establishes that at any step $t$, the following holds:
@@
If $t$ follows Case (a) of Lemma˜19, then
$h_{t}-h_{t+1}$ $\geq\frac{\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{d}_{t}\right\rangle^{2}}{2LD^{2}}$ by (17)
$\geq\frac{\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{v}_{t}^{A}-\mathbf{v}^{FW}_{t}\right\rangle^{2}}{8LD^{2}}$ by (15)
@@
If $t$ follows Case (b) of Lemma˜19, then:
@@
Denoting with $\mathcal{T}_{\mathrm{drop}}\subseteq[T]$ the set of drop steps, we can already observe that the sequence $\{h_{t}\}_{t\in[T]\backslash\mathcal{T}_{\mathrm{drop}}}$ follows the hypotheses of Lemma˜20, with parameters $c_{1},c_{2},\alpha$ defined in the statement of the theorem.
We can then observe that the number of drop steps is bounded by the number of FW steps on the sequence. This implies in the worst case that in $2t,t>1$ iterations, we can reach the primal gap of $t$ FW and pairwise steps combined. Let us denote $\tau$ the number of steps that were not drop steps up to iteration $t$, then:
@@
Lemma 18 (Geometric sharpness braun2022conditional, Lemma 3.32).
@@
Lemma 19 (Minimum decrease tsuji2021sparser, Lemma 3.4).
@@
Lemma 20 (Convergence rates from contraction inequalities, braun2022conditional, Lemma 2.21).
@@
Finally, we can exploit the recent results of zhao2025new on AFW for log-homogeneous self-concordant barrier functions to show that on D-Opt Problem, BPCG converges at a linear rate in primal and FW gaps.
Proposition 1 (Linear convergence of BPCG on D-Opt (zhao2025new)).
@@
The core observation required to extend zhao2025new is that the gap produced by BPCG upper-bounds the one produced by AFW. Let $\mathbf{d}^{\text{BPCG}},\mathbf{d}^{\text{AFW}}$ be the search directions produced by BPCG, AFW respectively. A drop step can occur in both algorithms in at most half of the iterations, since each drop step reduces the size of the active set by one. We can focus on “effective” steps in which no vertex is dropped. BPCG either takes a local pairwise step with direction:
@@
and $\mathcal{S}_{t}$ the local active set. Using the same notation, $\mathbf{d}^{\text{AFW}}=\mathbf{a}-\mathbf{x}$. The quantity determining whether a local pairwise step is selected is $\left\langle\nabla f(\mathbf{x}),\mathbf{d}^{\text{BPCG}}\right\rangle$, which needs to be greater than the FW gap $\left\langle\nabla f(\mathbf{x}),\mathbf{x}-\mathbf{v}\right\rangle$ at the current point. Regardless of the direction taken by BPCG, tsuji2021sparser provides:
@@
which is precisely the only inequality required from effective AFW steps. The rest of the proofs from zhao2025new follow without relying on the AFW descent direction specifically. ∎
Proposition˜1 unfortunately relies on the fact that the D-Opt objective is a log-barrier in addition to being self-concordant. Since this is not the case for the trace inverse function, we cannot transpose this result to the A-Opt Problem.